# Martingale Characterizations of Non-Homogeneous Counting Processes and Their Fractional Variants

### v1

@@ Abstract.
This paper investigates the martingale characterizations of non-homogeneous counting processes and their fractional generalizations. We show that the weighted sum of non-homogeneous Poisson processes (NPPs) is the non-homogeneous generalized counting process (NGCP). Both the compensated and exponential forms of martingale characterization for NGCP are obtained, and are shown to be equivalent. Moreover, we provide martingale characterizations for various time-changed variants of the NGCP and their Skellam versions using stable and/or inverse stable subordinators.
@@ 3.1. Martingale characterization of NPP
where the last step follows because $\{f(X(s))\}_{t\geq 0}$ has finitely many jumps. Now, as $\{f(X(t-))\}_{t\geq 0}$ is predictable and $\{\mathcal{N}_{j}(t)-\Lambda_{j}(t)\}_{t\geq 0}$ is a $\mathcal{F}_{t}^{j}$-martingale, by Theorem 11.4.5 of Shreve2004 $\{\exp\left(X(t)\right)\}_{t\geq 0}$ is a $\mathcal{F}_{t}$-martingale. ∎
@@ 3.3. Martingale characterization of NGCP
We begin with the following result that shows the unique representation of a NGCP as a weighted sum of NPPs.
@@ 4. Martingale characterizations of fractional variants
In this section, we investigate martingale characterizations for various time-changed variants of the processes considered in Section 3 and their Skellam versions. First note that the martingale characterization of a NGCP (see Theorems 3.7 and 3.8) assumes that the underlying NPPs are independent. However, for fractional variants of NGCP, independence among NPPs is not preserved after time changing them by a common stable and/or inverse stable subordinator. So a different approach is required to establish martingale characterizations in such cases. We obtain the compensated martingale characterizations for these fractional variants. The exponential martingale characterizations follow from their equivalence (see Theorems 3.3 and 3.5), which is independent of any fractional parameter. We begin with the following lemma (see Lemma 1 of Aletti2018) which will be used later.
@@
Let $\{K(t)\}_{t\geq 0}$ be a NTFPP so that $K(t)=\mathcal{N}\left(Y_{\alpha}(t)\right)$, where $\{Y_{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of the NPP $\{\mathcal{N}(t)\}_{t\geq 0}$ with rate function $\lambda(t)$.
For the NPP $\{\mathcal{N}(t)\}_{t\geq 0}$ the compensated process $\{\mathcal{N}(t)-\Lambda(t)\}_{t\geq 0}$ is a right-continuous martingale with respect to the natural filtration of $\mathcal{N}(t)$. Define
@@
First we show that $\{X(t)\}_{t\geq 0}$ is a martingale with respect to $\mathcal{F}_{t}=\sigma\left(K(s),s\leq t\right)\vee\sigma\left(Y_{\alpha}(s),s\geq 0\right)$.
For $0\leq s<t$, using the tower property and conditioning on $Y_{\alpha}(t)$, we have
@@
Note that $\mathcal{N}(Y_{\alpha}(t))|Y_{\alpha}(t)\sim\text{Poi}(\Lambda(Y_{\alpha}(t)))$. Using the martingale property of a Poisson process, we get
@@
The martingale characterization of a non-homogeneous tempered space fractional Poisson process (NTSFPP) given below follows from Theorem 4.2 by taking $\alpha=1$ and observing that $Y_{1}(t)=t$ a.s.
Corollary 4.1.
Let $\{G(t)\}_{t\geq 0}$ be a simple locally finite point process. Then $\{G(t)\}_{t\geq 0}$ is a NTSFPP if and only if there exist a rate function $\lambda:[0,\infty)\to[0,\infty)$ with $\Lambda(t)=\int_{0}^{t}\lambda(u)du<\infty$ and a tempered stable subordinator $\{D_{\beta,\theta}(t)\}_{t\geq 0}$ such that the process
$\{H(t)\}_{t\geq 0}=\{G(t)-\Lambda(D_{\beta,\theta}(t))\}_{t\geq 0}$
is a right-continuous martingale with respect to the induced filtration $\mathcal{F}_{t}=\sigma\left(G(s),s\leq t\right)\vee\sigma\left(D_{\beta,\theta}(s),s\geq 0\right)$ and for any $T>0$,
$\{H(\tau),\,\tau\,\,\text{is stopping time such that}\,\,D_{\beta,\theta}(\tau)\leq T\}$
is uniformly integrable.
Next, we investigate the martingale characterizations of the generalized versions of NTFPP, NTSFPP and NTSTFPP which allow multiple arrivals with different rate functions. The following result gives the martingale characterization of a non-homogeneous generalized fractional counting process (NGFCP), the generalized version of NTFPP, defined as
@@
Let $\{\mathcal{N}_{1}(t)\}_{t\geq 0}$, $\{\mathcal{N}_{2}(t)\}_{t\geq 0}$$\dots\{\mathcal{N}_{k}(t)\}_{t\geq 0}$ be $k$ independent NPPs with respective rate functions $\lambda_{1}(t)$, $\lambda_{2}(t)$, $\dots,~\lambda_{k}(t)$. Let $\{Y_{\alpha}(t)\}_{t\geq 0}$ be an inverse stable subordinator independent of these NPPs. Then, by Theorem 3.6, we have $\mathcal{M}(t)=\sum_{j=1}^{k}j\mathcal{N}_{j}(t)$ which implies $\mathcal{M}(Y_{\alpha}(t))=\sum_{j=1}^{k}j\mathcal{N}_{j}(Y_{\alpha}(t))$, that is, $\mathcal{M}^{\alpha}(t)=\sum_{j=1}^{k}j\mathcal{N}_{j}^{\alpha}(t)$ where $\{\mathcal{N}_{j}^{\alpha}(t)\}_{t\geq 0}$, $1\leq j\leq k$ are NTFPPs. Theorem 4.1 implies that $\{X_{j}(t)=\mathcal{N}_{j}^{\alpha}(t)-\Lambda_{j}(Y_{\alpha}(t))\}_{\tau\geq 0}$ is a $\{\mathcal{F}_{t}^{j}\}_{t\geq 0}$-martingale and that for any $T>0$, the family $\{X_{j}(\tau),\,\tau\,\,$ is stopping time such
@@
Corollary 4.3.
Let $\{\mathcal{M}^{\alpha}_{\beta,\theta}(t)\}_{t\geq 0}$ be a NTGSTFCP. Then the process $\{\mathcal{M}^{\alpha}_{\beta,\theta}(t)-\sum_{j=1}^{k}j\Lambda_{j}\left(D_{\beta,\theta}(Y_{\alpha}(t))\right)\}_{t\geq 0}$ is a $\{\mathcal{F}_{t}\}_{t\geq 0}$-martingale, where $\Lambda_{j}(t)=\int_{0}^{t}\lambda_{j}(u)du<\infty$ for some rate function $\lambda:[0,\infty)\to[0,\infty)$, $\{D_{\beta,\theta}(t)\}_{t\geq 0}$ is a tempered stable subordinator, $\{Y_{\alpha}(t)\}_{t\geq 0}$ is an independent inverse stable subordinator and
$\mathcal{F}_{t}=\sigma\left(\mathcal{N}^{\alpha}_{1\beta,\theta}(s),~\mathcal{N}^{\alpha}_{2\beta,\theta}(s),\dots,~\mathcal{N}^{\alpha}_{k\beta,\theta}(s),~s\leq t\right)\vee\sigma\left(D_{\beta,\theta}(Y_{\alpha}(s)),s\geq 0\right)$.
Now we provide the martingale characterization of a non-homogeneous tempered generalized space fractional counting process (NTGSFCP), the generalized version of NTSFPP, defined as
@@
where $\{\mathcal{M}(t)\}_{t\geq 0}$ is a NGCP and $\{D_{\beta,\theta}(t)\}_{t\geq 0}$ is an independent tempered stable subordinator.
Theorem 4.5.
The process $\{\mathcal{M}_{\beta,\theta}(t)\}_{t\geq 0}$ is a NTGSFCP if and only if there exist $k$ simple locally finite point processes $\{\mathcal{N}_{j\beta,\theta}(t)\}_{t\geq 0}$, $1\leq j\leq k$ such that $\mathcal{M}_{\beta,\theta}(t)=\sum_{j=1}^{k}j\mathcal{N}_{j\beta,\theta}(t)$ and $X_{j}(t)=\mathcal{N}_{j\beta,\theta}(t)-\Lambda_{j}(D_{\beta,\theta}(t))$, $t\geq 0$ is a right continuous $\{\mathcal{F}_{t}^{j}\}_{t\geq 0}$-martingale, where $\Lambda_{j}(t)=\int_{0}^{t}\lambda_{j}(u)du<\infty$ for some rate function $\lambda_{j}:[0,\infty)\to[0,\infty)$, $\{D_{\beta,\theta}(t)\}_{t\geq 0}$ is a tempered stable subordinator and $\mathcal{F}_{t}^{j}=\sigma\left(\mathcal{N}_{\beta,\theta_{j}}(s),s\leq t\right)\vee\sigma\left(D_{\beta,\theta}(s),s\geq 0\right)$. Moreover, for any $T>0$, $\{X_{j}(\tau),\,\tau\,\,\text{is stopping time such that}\,\,D_{\beta,\theta}(\tau)\leq T\}$ is uniformly integrable.
Proof.
The proof follows from Theorem 4.4 by taking $\alpha=1$ and noting that $Y_{1}(t)=t$ a.s.
The result below follows from Theorem 4.5 and is similar to that of Corollary 4.3 with $\alpha=1$.
Corollary 4.4.
Let $\{\mathcal{M}_{\beta,\theta}(t)\}_{t\geq 0}$ be a NTGSFCP. Then the process $\{\mathcal{M}_{\beta,\theta}(t)-\sum_{j=1}^{k}j\Lambda_{j}\left(D_{\beta,\theta}(t)\right)\}_{t\geq 0}$ is a $\{\mathcal{F}_{t}\}_{t\geq 0}$-martingale, where $\Lambda_{j}(t)=\int_{0}^{t}\lambda_{j}(u)du<\infty$ for some rate function $\lambda:[0,\infty)\to[0,\infty)$, $\{D_{\beta,\theta}(t)\}_{t\geq 0}$ is a tempered stable subordinator and $\mathcal{F}_{t}=\sigma\left(\mathcal{N}_{1\beta,\theta}(s),~\mathcal{N}_{2\beta,\theta}(s),\dots,~\mathcal{N}_{k\beta,\theta}(s),~s\leq t\right)\vee\sigma\left(D_{\beta,\theta}(s),s\geq 0\right)$.
Next, we investigate the martingale characterization of the mixed fractional extension of a NGCP. The non-homogeneous mixed fractional Poisson process (NMFPP) is defined as
@@
Corollary 4.5.
Let $\{\mathcal{M}^{\alpha_{1},\alpha_{2}}(t)\}_{t\geq 0}$ be a NMFCP. Then the process $\{\mathcal{M}^{\alpha_{1},\alpha_{2}}(t)-\sum_{j=1}^{k}j\Lambda_{j}\left(Y_{\alpha_{1},\alpha_{2}}(t)\right)\}_{t\geq 0}$ is a $\{\mathcal{F}_{t}\}_{t\geq 0}$-martingale, where $\Lambda_{j}(t)=\int_{0}^{t}\lambda_{j}(u)du<\infty$ for some rate function $\lambda:[0,\infty)\to[0,\infty)$, $\{Y_{\alpha_{1},\alpha_{2}}(t)\}_{t\geq 0}$ is an mixed inverse stable subordinator and $\mathcal{F}_{t}=\sigma\left(\mathcal{N}_{1}^{\alpha_{1},\alpha_{2}}(s),~\mathcal{N}_{2}^{\alpha_{1},\alpha_{2}}(s),\dots,~\mathcal{N}_{k}^{\alpha_{1},\alpha_{2}}(s),~s\leq t\right)\\ \vee\sigma\left(Y_{\alpha_{1},\alpha_{2}}(s),s\geq 0\right)$.
Remark 4.3.
For $k=1$, Theorem 4.6 gives the martingale characterization of a NMFPP.
Finally, we obtain the martingale characterizations of Skellam processes. In general, a Skellam process is defined as the difference between two idependent counting processes. The non-homogeneous generalized Skellam process (NGSP) is defined as the difference between two independent NGCPs. Specifically, let $\{\mathcal{M}_{1}(t)\}_{t\geq 0}$ and $\{\mathcal{M}_{2}(t)\}_{t\geq 0}$ be two independent NGCPs with respective intensity functions $\lambda_{j}:[0,\infty)\rightarrow[0,\infty)$ and $\mu_{j}:[0,\infty)\rightarrow[0,\infty)$ for $j=1,2,\ldots,k$, and with initial conditions $\mathcal{M}_{1}(0)=\mathcal{M}_{2}(0)=0$. Then the NGSP is given by $\{\mathcalboondox{S}(t)=\mathcal{M}_{1}(t)-\mathcal{M}_{2}(t)\}_{t\geq 0}$ whose martingale characterization is given as follows.
@@
The following result provides the Watanabe characterization of a non-homogeneous time fractional Skellam process (NFSP) which is obtained by time changing a non-homogeneous Skellam process (NSP) with an independent inverse stable subordinator. It can also be obtained as the difference of two independent NTFPPs.
@@
Let $\{\mathcal{K}(t)\}_{t\geq 0}$ be a simple locally finite point process. Then, $\{\mathcal{K}(t)\}_{t\geq 0}$ is a NFSP if and only if there exist rate functions $\lambda:[0,\infty)\to[0,\infty)$ and $\mu:[0,\infty)\to[0,\infty)$ with $\Lambda(t)=\int_{0}^{t}\lambda(u)du<\infty$ and $T(t)=\int_{o}^{t}\mu(u)du<\infty$ and an independent inverse stable subordinator $\{Y_{\alpha}(t)\}_{t\geq 0}$ such that the process
$\{\mathcal{X}(t)\}_{t\geq 0}=\{\mathcal{K}(t)-\left(\Lambda(Y_{\alpha}(t))-T(Y_{\alpha}(t))\right)\}_{t\geq 0}$ (4.28)
is a right-continuous martingale with respect to the induced filtration $\mathcal{F}_{t}=\sigma\left(\mathcal{K}(s),s\leq t\right)\vee\sigma\left(Y_{\alpha}(s),s\geq 0\right)$ and for any $T>0$,
@@
Let $\mathcal{K}(t)=\mathcal{N}_{1}(Y_{\alpha}(t))-\mathcal{N}_{2}(Y_{\alpha}(t))$, where $\{\mathcal{N}_{1}(Y_{\alpha}(t))\}_{t\geq 0}$ and $\{\mathcal{N}_{2}(Y_{\alpha}(t))\}_{t\geq 0}$ are independent NTFPPs with parameters $\Lambda(t)$ and $T(t)$, and $\{Y_{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of these processes. Then the process $\{\mathcal{K}(t)\}_{t\geq 0}$ represents a non-homogeneous fractional Skellam process (NFSP).
By Theorem 4.1, $\{\mathcal{N}_{1}(Y_{\alpha}(t))-\Lambda(Y_{\alpha}(t))\}_{t\geq 0}$ and $\{\mathcal{N}_{2}(Y_{\alpha}(t))-T(Y_{\alpha}(t))\}_{t\geq 0}$ are $\mathcal{F}_{t}$-martingales. This implies that $\{\mathcal{X}(t)\}_{t\geq 0}$ is a $\mathcal{F}_{t}$-martingale since the difference of two martingales is again a martingale with respect to a common filtration. Now consider the processes $\{\mathcal{N}_{1}(Y_{\alpha}(\tau))-\Lambda(Y_{\alpha}(\tau))\}$ and $\{\mathcal{N}_{2}(Y_{\alpha}(\tau))-T(Y_{\alpha}(\tau))\}$ where $\tau$ is a stopping time such that $Y_{\alpha}(\tau)\leq T$. By Theorem 4.1, these processes are uniformly integrable. Since uniform integrability is closed under linear combinations, the process given by (4.29) is also uniformly integrable.
The converse part follows from the converse part in the proof of Theorem 4.1 applied to the two NTFPPs, and then taking their difference. ∎
The NGFSP is a process obtained by time changing a NGSP with an independent inverse stable subordinator or by taking the difference between two independent NGFCPs. The martingale characterization of a NGFSP is given below.
@@
The process $\{\mathcalboondox{S}^{\alpha}(t)\}_{t\geq 0}$ is a NGFSP if and only if there exist $k$ simple locally finite point processes $\{\mathcal{S}^{\alpha}_{j}(t)\}_{t\geq 0}$, $1\leq j\leq k$ such that $\mathcalboondox{S}^{\alpha}(t)=\sum_{j=1}^{k}j\mathcal{S}^{\alpha}_{j}(t)$ and $X_{j}(t)=\mathcal{S}^{\alpha}_{j}(t)-(\Lambda_{j}(Y_{\alpha}(t))-T_{j}(Y_{\alpha}(t))$, $t\geq 0$ is a right continuous $\{\mathcal{F}_{t}^{j}\}_{t\geq 0}$-martingale, where $\mathcal{F}_{t}^{j}=\sigma\left(\mathcal{S}_{j}^{\alpha}(s),s\leq t\right)\vee\sigma\left((Y_{\alpha}(s),s\geq 0\right)$ and $\{Y_{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of $\{S_{j}(t)\}_{t\geq 0}$. Moreover, for any $T>0$, $\{X_{j}(\tau),\\ \tau\,\,\text{is stopping time such that}\,\,(Y_{\alpha}(\tau)\leq T\}$ is uniformly integrable.

### v2

@@ Abstract.
This paper investigates the martingale characterizations of non-homogeneous counting processes and their fractional generalizations. We show that a weighted sum of non-homogeneous Poisson processes (NPPs) is the non-homogeneous generalized counting process (NGCP). Both the compensated and exponential forms of martingale characterization for NGCP are obtained, and are shown to be equivalent. Moreover, we provide martingale characterizations for various time-changed variants of the NGCP and their Skellam versions using stable and/or inverse stable subordinators.
@@ 3.1. Martingale characterization of NPP
where the last step follows because $\{f(X(s))\}_{t\geq 0}$ has finitely many jumps. Now, as $\{f(X(t-))\}_{t\geq 0}$ is predictable and $\{\mathcal{N}_{j}(t)-\Lambda_{j}(t)\}_{t\geq 0}$ is a $\mathcal{F}_{t}^{j}$-martingale, by Theorem 11.4.5 of Shreve2004, $\{\exp\left(X(t)\right)\}_{t\geq 0}$ is a $\mathcal{F}_{t}$-martingale. ∎
@@ 3.3. Martingale characterization of NGCP
We begin with the following result that shows the unique representation of a NGCP as a weighted sum of NPPs where the weights are the jump sizes of the NGCP.
@@ 4. Martingale characterizations of fractional variants
Tathe2025 obtained martingale characterizations for some fractional homogeneous processes. In this section, we investigate martingale characterizations for various time-changed variants of the non-homogeneous processes considered in Section 3 and their Skellam versions. First, note that the martingale characterization of a NGCP (see Theorems 3.7 and 3.8) assumes that the underlying NPPs are independent. However, for fractional variants of NGCP, independence among NPPs is not preserved after time changing them by a common stable and/or inverse stable subordinator. So a different approach is required to establish martingale characterizations in such cases. We obtain the compensated martingale characterizations for these fractional variants. The exponential martingale characterizations follow from their equivalence (see Theorems 3.3 and 3.5), which is independent of any fractional parameter. We begin with the following lemma (see Lemma 1 of Aletti2018) which will be used later.
@@
Let $\{K(t)\}_{t\geq 0}$ be a NTFPP so that $K(t)=\mathcal{N}\left(Y_{\alpha}(t)\right)$, where $\{Y_{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of the NPP $\{\mathcal{N}(t)\}_{t\geq 0}$ with rate function $\lambda(t)$. Then
@@
First we show that $\{X(t)\}_{t\geq 0}$ is a martingale with respect to $\mathcal{F}_{t}=\sigma\left(K(s),s\leq t\right)\vee\sigma\left(Y_{\alpha}(s),s\geq 0\right)$. For $0\leq s<t$, using the tower property and conditioning on $Y_{\alpha}(t)$, we have
@@
Note that $\mathcal{N}(Y_{\alpha}(t))|Y_{\alpha}(t)\sim\text{Poi}(\Lambda(Y_{\alpha}(t)))$. Using the martingale property of a non-homogeneous Poisson process (see Theorem 3.2), we get
@@
The martingale characterization of a NTSFPP follows from Theorem 4.2 by taking $\alpha=1$ and observing that $Y_{1}(t)=t$ a.s. Next, we investigate martingale characterizations of the generalized versions of NTFPP, NTSFPP and NTSTFPP which allow multiple arrivals with different rate functions. The next result gives the martingale characterization of a non-homogeneous generalized fractional counting process (NGFCP), the generalized version of NTFPP, defined as
@@
Let $\{\mathcal{N}_{1}(t)\}_{t\geq 0}$, $\{\mathcal{N}_{2}(t)\}_{t\geq 0}$$\dots,\{\mathcal{N}_{k}(t)\}_{t\geq 0}$ be $k$ independent NPPs with respective rate functions $\lambda_{1}(t)$, $\lambda_{2}(t)$, $\dots,~\lambda_{k}(t)$. Let $\{Y_{\alpha}(t)\}_{t\geq 0}$ be an inverse stable subordinator independent of these NPPs. Then, by Theorem 3.6, we have $\mathcal{M}(t)=\sum_{j=1}^{k}j\mathcal{N}_{j}(t)$ which implies $\mathcal{M}(Y_{\alpha}(t))=\sum_{j=1}^{k}j\mathcal{N}_{j}(Y_{\alpha}(t))$, that is, $\mathcal{M}^{\alpha}(t)=\sum_{j=1}^{k}j\mathcal{N}_{j}^{\alpha}(t)$ where $\{\mathcal{N}_{j}^{\alpha}(t)\}_{t\geq 0}$, $1\leq j\leq k$ are NTFPPs. Theorem 4.1 implies that $\{X_{j}(t)=\mathcal{N}_{j}^{\alpha}(t)-\Lambda_{j}(Y_{\alpha}(t))\}_{\tau\geq 0}$ is a $\{\mathcal{F}_{t}^{j}\}_{t\geq 0}$-martingale and that for any $T>0$, the family $\{X_{j}(\tau),\,\tau\,\,$ is stopping time such
@@
A non-homogeneous tempered generalized space fractional counting process (NTGSFCP), the generalized version of NTSFPP, is defined as
@@
where $\{\mathcal{M}(t)\}_{t\geq 0}$ is a NGCP and $\{D_{\beta,\theta}(t)\}_{t\geq 0}$ is an independent tempered stable subordinator. The martingale characterization of NTGSFCP follows from Theorem 4.4 by taking $\alpha=1$ and noting that $Y_{1}(t)=t$ a.s. Next, we investigate the martingale characterization of the mixed fractional extension of a NGCP. The non-homogeneous mixed fractional Poisson process (NMFPP) is defined as
@@
Finally, we obtain the martingale characterizations of Skellam processes. In general, a Skellam process is defined as the difference of two idependent counting processes. The non-homogeneous generalized Skellam process (NGSP) is defined as the difference of two independent NGCPs. Specifically, let $\{\mathcal{M}_{1}(t)\}_{t\geq 0}$ and $\{\mathcal{M}_{2}(t)\}_{t\geq 0}$ be two independent NGCPs with respective intensity functions $\lambda_{j}:[0,\infty)\rightarrow[0,\infty)$ and $\mu_{j}:[0,\infty)\rightarrow[0,\infty)$ for $j=1,2,\ldots,k$, and with initial conditions $\mathcal{M}_{1}(0)=\mathcal{M}_{2}(0)=0$. Then the NGSP is given by $\{\mathcalboondox{S}(t)=\mathcal{M}_{1}(t)-\mathcal{M}_{2}(t)\}_{t\geq 0}$ whose martingale characterization is given as follows.
@@
The following result provides the Watanabe characterization of a non-homogeneous time fractional Skellam process (NFSP) which may be defined as the difference of two independent NTFPPs.
@@
Let $\{\mathcal{K}_{1}(t)\}_{t\geq 0}$ be a simple locally finite point process. Then, $\{\mathcal{K}_{1}(t)\}_{t\geq 0}$ is a NFSP if and only if there exist rate functions $\lambda:[0,\infty)\to[0,\infty)$ and $\mu:[0,\infty)\to[0,\infty)$ with $\Lambda(t)=\int_{0}^{t}\lambda(u)du<\infty$ and $T(t)=\int_{o}^{t}\mu(u)du<\infty$ and two independent inverse stable subordinators $\{Y_{\alpha_{1}}(t)\}_{t\geq 0}$ and $\{Y_{\alpha_{2}}(t)\}_{t\geq 0}$ such that the process
$\{\mathcal{X}(t)\}_{t\geq 0}=\{\mathcal{K}_{1}(t)-\left(\Lambda(Y_{\alpha_{1}}(t))-T(Y_{\alpha_{2}}(t))\right)\}_{t\geq 0}$ (4.28)
is a right-continuous martingale with respect to the induced filtration $\mathcal{F}_{t}=\sigma\left(\mathcal{K}_{1}(s),s\leq t\right)\vee\sigma\left(Y_{\alpha_{1}}(s),s\geq 0\right)\vee\sigma\left(Y_{\alpha_{2}}(s),s\geq 0\right)$ and for any $T>0$,
$\{\mathcal{X}(\tau),\,\tau\,\,\text{is stopping time such that}\,\,\max\left(Y_{\alpha_{1}}(\tau),Y_{\alpha_{2}}(\tau)\right)\leq T\}$ (4.29)
is uniformly integrable.
Proof.
Assume $\{\mathcal{K}_{1}(t)\}_{t\geq 0}$ is a NFSP. By definition, $\mathcal{K}_{1}(t)=\mathcal{N}_{1}(Y_{\alpha_{1}}(t))-\mathcal{N}_{2}(Y_{\alpha_{2}}(t))$, where $\{\mathcal{N}_{1}(Y_{\alpha_{1}}(t))\}_{t\geq 0}$ and $\{\mathcal{N}_{2}(Y_{\alpha_{2}}(t))\}_{t\geq 0}$ are independent NTFPPs with parameters $\Lambda(t)$ and $T(t)$, and $\{Y_{\alpha_{1}}(t)\}_{t\geq 0}$ and $\{Y_{\alpha_{2}}(t)\}_{t\geq 0}$ are two independent inverse stable subordinators that are also independent of these processes.
By Theorem 4.1, $\{\mathcal{N}_{1}(Y_{\alpha_{1}}(t))-\Lambda(Y_{\alpha_{1}}(t))\}_{t\geq 0}$ and $\{\mathcal{N}_{2}(Y_{\alpha_{2}}(t))-T(Y_{\alpha_{2}}(t))\}_{t\geq 0}$ are $\mathcal{F}_{t}$-martingales. This implies that $\{\mathcal{X}(t)\}_{t\geq 0}$ is a $\mathcal{F}_{t}$-martingale since the difference of two martingales with respect to a common filtration is a martingale. Now consider the processes $\{\mathcal{N}_{1}(Y_{\alpha_{1}}(\tau))-\Lambda(Y_{\alpha_{1}}(\tau))\}$ and $\{\mathcal{N}_{2}(Y_{\alpha_{2}}(\tau))-T(Y_{\alpha_{2}}(\tau))\}$ where $\tau$ is a stopping time such that $\max\left(Y_{\alpha_{1}}(\tau),~Y_{\alpha_{2}}(\tau)\right)\leq T$. By Theorem 4.1, these processes are uniformly integrable. Since uniform integrability is closed under linear combinations, the process given by (4.29) is also uniformly integrable.
The converse part follows from the converse part in the proof of Theorem 4.1 applied to the two NTFPPs $\{\mathcal{N}_{1}(Y_{\alpha_{1}}(t))\}_{t\geq 0}$ and $\{\mathcal{N}_{2}(Y_{\alpha_{2}}(t))\}_{t\geq 0}$, and then taking their difference. ∎
If a NFSP is defined by time-changing a non-homogeneous Skellam process (NSP) with an independent inverse stable subordinator, the following result provides its Watanabe characterization.
Theorem 4.8.
Let $\{\mathcal{K}_{2}(t)\}_{t\geq 0}$ be a simple locally finite point process. Then, $\{\mathcal{K}_{2}(t)\}_{t\geq 0}$ is a NFSP if and only if there exist rate functions $\lambda:[0,\infty)\to[0,\infty)$ and $\mu:[0,\infty)\to[0,\infty)$ with $\Lambda(t)=\int_{0}^{t}\lambda(u)du<\infty$ and $T(t)=\int_{o}^{t}\mu(u)du<\infty$ and an independent inverse stable subordinator $\{Y_{\alpha}(t)\}_{t\geq 0}$ such that the process
$\{\mathcal{X}(t)\}_{t\geq 0}=\{\mathcal{K}_{2}(t)-\left(\Lambda(Y_{\alpha}(t))-T(Y_{\alpha}(t))\right)\}_{t\geq 0}$ (4.30)
is a right-continuous martingale with respect to the induced filtration $\mathcal{F}_{t}=\sigma\left(\mathcal{K}_{2}(s),s\leq t\right)\vee\sigma\left(Y_{\alpha}(s),s\geq 0\right)$ and for any $T>0$,
@@
For a non-homogeneous Skellam process $\{\mathcal{S}(t)\}_{t\geq 0}$, note that $\mathcal{K}_{2}(t)\big|Y_{\alpha}(t)=\mathcal{S}\left(Y_{\alpha}(t)\right)\big|Y_{\alpha}(t)$ has a non-homogeneous Skellam distribution with mean $\Lambda(Y_{\alpha}(t))-T(Y_{\alpha}(t))$. It can also be shown that $\mathbb{E}(X(\tau))=0$ and $\mathbb{E}\left(X^{2}(\tau)\right)\leq\Lambda(T)+T(T)$. The proof of the result follows in a manner analogous to the proof of Theorem 4.1. ∎
The following result establishes the Watanabe characterization of NGFSP obtained by time-changing a NGSP with an independent inverse stable subordinator.
@@
The process $\{\mathcalboondox{S}^{\alpha}(t)\}_{t\geq 0}$ is a NGFSP if and only if there exist $k$ simple locally finite point processes $\{\mathcal{S}^{\alpha}_{j}(t)\}_{t\geq 0}$, $1\leq j\leq k$ such that $\mathcalboondox{S}^{\alpha}(t)=\sum_{j=1}^{k}j\mathcal{S}^{\alpha}_{j}(t)$ and $X_{j}(t)=\mathcal{S}^{\alpha}_{j}(t)-(\Lambda_{j}(Y_{\alpha}(t))-T_{j}(Y_{\alpha}(t))$, $t\geq 0$ is a right-continuous $\{\mathcal{F}_{t}^{j}\}_{t\geq 0}$-martingale, where $\mathcal{F}_{t}^{j}=\sigma\left(\mathcal{S}_{j}^{\alpha}(s),s\leq t\right)\vee\sigma\left(Y_{\alpha}(s),s\geq 0\right)$ and $\{Y_{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of $\{\mathcal{S}_{j}(t)\}_{t\geq 0}$. Moreover, for any $T>0$, $\{X_{j}(\tau),\tau\,\,\text{is stopping time such that}\,\,(Y_{\alpha}(\tau))\leq T\}$ is uniformly integrable.